# Reducing Energy Bloat in Large Model Training

### v2

@@ Abstract.
Training large AI models on numerous GPUs consumes a massive amount of energy, making power delivery one of the largest limiting factors in building and operating datacenters for AI workloads. However, we observe that not all energy consumed during training directly contributes to end-to-end throughput, and a significant portion can be removed without slowing down training, which we call energy bloat.
In this work, we identify two independent sources of energy bloat in large model training and propose Perseus, a training system that mitigates both. To do this, Perseus obtains the “iteration time–energy” Pareto frontier of any large model training job using an efficient graph cut-based algorithm and schedules the energy consumption of computations across time to remove both types of energy bloat. Evaluation on large models including GPT-3 and Bloom shows that Perseus reduces the energy consumption of large model training by up to 30% without any throughput loss or hardware modification, enabling energy reduction – and therefore cost savings – otherwise unattainable before.
@@ 1. Introduction
As deep neural networks (DNNs) continue to grow in model and dataset size (scalinglaws-arxiv20, ; chinchilla-neurips22, ), the energy consumption of large model training is increasing as well. For instance, training GPT-3 (gpt3, ) reportedly consumed 1.3 GWh (patterson2021carbon, ). Then, this was recently dwarfed by a 200B model trained by Amazon consuming about 11.9 GWh (hamilton2024constraint, ), which is enough to power more than 1,000 average US households for a year (us-household, ). Such energy-intensive large model training workloads not only inflate datacenter operational expenses, but also made power delivery a primary challenge in building datacenters today (cbre2023, ; mckinsey2023, ; cbre2024, ).
@@
Despite recent works on accelerating large model training (megatronlm-sc21, ; alpa-osdi22, ; merak-tpds23, ; oobleck-sosp23, ), energy optimization still remains an open challenge (greenai-cacm20, ; patterson2021carbon, ). While energy optimization is well-studied in the hardware community (eyeriss-jssc16, ; vivienne17efficient, ; puma-asplos19, ; kws-jssc24, ), the power bottleneck of recent datacenters (cbre2023, ; mckinsey2023, ; cbre2024, ) shows that efficiency gains from hardware advancement alone are not sufficient to sustain the growing demand for AI compute. As such, a recent line of work demonstrated that software plays a significant role in energy optimization by capturing application characteristics that general-purpose hardware cannot (e.g., no need to finish computation before the deadline), bringing hardware-agnostic energy-efficiency gains (gpoeo, ; zeus-nsdi23, ; chase-ccai23, ; enemy-hpca23, ; envpipe-atc23, ).
@@
In this paper, we seek a software method that reduces the energy consumption of large model training without slowdown, thereby also reducing the average power draw. To that end, we identify energy bloat – the portion of energy consumption that can be removed without slowdown – in software systems for large model training. We find two independent sources of energy bloat: intrinsic and extrinsic, and propose a single optimization framework that minimizes both.
@@
Intrinsic energy bloat comes from computation imbalance when a large model is distributed across multiple GPUs with pipeline parallelism (§2.2). Balancing the amount of computation in each pipeline stage is an important problem for distributed execution planning (pipedream-sosp19, ; dapple-ppopp21, ; autopipe-icml21, ; alpa-osdi22, ), but perfectly balancing every stage is not always possible because layers in a DNN are coarse-grained tensor operations with varying amounts of computation. When all stages do not have the same computation time, those not on the critical path of computation run needlessly fast – i.e., they consume energy that does not contribute to the overall training throughput. Such intrinsic energy bloat opens up the opportunity to precisely slow down each non-critical computation in the pipeline such that the length of the critical path does not change.
@@
Extrinsic energy bloat, in contrast, arises when multiple pipelines run in parallel in a synchronous fashion to scale out training to massive datasets, and one or more pipelines run slower than the others (§2.3). Root causes behind such slowdowns are varied, including power/thermal throttling (A100powerthermal-energies21, ; gpupower-cal23, ; thunderbolt-osdi20, ; mvpp-asplos20, ; polca-arxiv23, ), I/O bottlenecks in the storage/network (datastalls-vldb21, ; recsys-dsi-isca22, ), hardware/software failures (bamboo-nsdi23, ; oobleck-sosp23, ), etc., and the likelihood of their presence increases with the scale and duration of training (largescalefailures-sc17, ; philly-atc19, ; mlaas-nsdi22, ). All pipelines running faster than the straggler pipeline are needlessly fast and wasting energy that does not affect the overall training throughput. Thus, extrinsic energy bloat allows us to slow down entire pipelines (while still keeping their intrinsic energy bloat low) without delaying gradient synchronization.
@@
In this work, we propose Perseus, which formulates a unified optimization framework to remove both intrinsic and extrinsic energy bloat from large model training (§3). At its core, Perseus efficiently pre-characterizes the entire “iteration time–energy” Pareto frontier, allowing it to minimize intrinsic bloat under normal operation and to mitigate extrinsic bloat arising from stragglers. Existing works fall short on both fronts. EnvPipe (envpipe-atc23, ) is limited to intrinsic bloat reduction with a point solution that leads to suboptimal energy reduction. Zeus (zeus-nsdi23, ), in contrast, ignores intrinsic bloat as it only considers single-GPU training, which also renders its “iteration time–energy” frontier suboptimal for large models.
@@
We show that characterizing the “iteration time–energy” Pareto frontier is NP-hard not only to solve but also to approximate within any constant factor. Given this impasse, we propose an efficient algorithm that optimally solves a relaxed problem instead (§4). At its core, Perseus represents one training iteration as a directed acyclic graph (DAG). Each node in the DAG represents either a forward or backward computation in a particular stage, and nodes are connected with edges that represent computation dependencies. Perseus’s optimization algorithm then efficiently generates all energy schedules, defined as the planned time and energy consumption of each DAG node (computation), that are on the “iteration time–energy” Pareto frontier. In doing so, instead of finding each point on the frontier individually, Perseus employs a graph cut-based algorithm that iteratively crawls up the frontier from the bottom, generating the next Pareto-optimal energy schedule by carefully tweaking its neighbor energy schedule on the frontier. Minimizing intrinsic and/or extrinsic energy bloat is then as simple as choosing the appropriate energy schedule from the Pareto frontier.
@@
Perseus is split into a server and a client library (§5). The client library can be integrated into any large model training framework as well as any accelerator that has an energy measurement and execution speed control interface, giving the server an abstract view of training computations and their time and energy consumption. The server produces Pareto-optimal energy schedules, which is realized by the client.
@@
Evaluation on large models including GPT-3 (gpt3, ), BERT (bert, ), T5 (t5-jmlr20, ), Bloom (bloom-arxiv22, ), and scaled-up Wide-ResNet (wideresnet-bmvc16, ) shows that Perseus is able to reduce per-iteration energy consumption by up to 30% with negligible or no slowdown, reducing both energy consumption and average power draw (§6).
@@
  • We propose Perseus, a software-only energy optimization system that removes energy bloat through a unified optimization framework and a graph cut-based algorithm.
@@ 2.1. Large Model Training
Large model training today is mostly dominated by 3D (data, tensor, and pipeline) parallelism (megatronlm-sc21, ; megatron-turing-arxiv22, ; bloom-arxiv22, ; merak-tpds23, ; falcon-arxiv23, ; llama3-blog, ; deepspeed-github, ; megatronlm-github, ; gpt-neox-github, ). Especially, pipeline parallelism partitions a large model and its training batch into multiple stages and microbatches respectively, and pipelines forward and backward computations through stages. Then, such pipelines are replicated multiple times to perform data parallel training. Since all pipelines must synchronize gradients at the end of each iteration, computations for the next iteration can only begin when all pipelines have finished and exchanged gradients.
@@ 2.2. Intrinsic Energy Bloat
Figure 1. One training iteration of GPT-3 1.3B with 4 pipeline stages and 6 microbatches on NVIDIA A100 GPUs drawn to scale. For example, F5 and B5 in the S2 row denote Forward and Backward for the fifth microbatch on Stage 2. The critical path of computation is traced with a blue line. Colors show power consumption. Especially, the blue background indicates that GPUs consume power when blocking on communication. Other models are visualized in Appendix A.
@@
Table 1. Latency ratio of the longest to the shortest stage on A100 GPUs. 1.00 would mean perfect balance.
We profile GPT-3 1.3B on NVIDIA A100 GPUs and visualize the timeline of one training iteration in Figure 1(a). In addition to the familiar bubbles in the 1F1B schedule (megatronlm-sc21, ), we observe gaps in between forward and backward computations, where the GPU is simply blocking on communication with an adjacent stage. Such gaps exist because the computation time of each pipeline stage is not perfectly balanced. Partitioning stages in a balanced manner is an important problem in distributed execution planning (gpipe-neurips19, ; pipedream-sosp19, ; dapple-ppopp21, ; alpa-osdi22, ), but perfect balancing is difficult because DNNs are essentially a sequence of coarse-grained tensor operations with varying size.
@@
Given stage imbalance, as we have seen in Figure 1(a), not all forward and backward computations are on the critical path of computation. This means that non-critical computations running at their maximum speed is not contributing to faster iteration time, and thus simply wasting energy. We call this intrinsic energy bloat, which can be reduced by precisely slowing down each non-critical computation without lengthening the critical path (Figure 1(b)). Although seemingly simple, this problem is not only NP-hard to solve, but also NP-hard to even approximate to any constant factor (svensson2012hardness, ).
@@ 2.3. Extrinsic Energy Bloat
Numerous replicas of the same pipeline run in a data parallel fashion in large model training. Because every pipeline must synchronize gradients at the end, if a pipeline runs slower, all other pipelines must wait until the straggler pipeline finishes (Figure 2(a)). Since the straggler pipeline determines end-to-end iteration time, all other pipelines running at their fastest possible iteration time is wasteful. We call this extrinsic energy bloat, because unlike intrinsic energy bloat, its cause is extrinsic to the pipeline. To reduce extrinsic bloat while keeping intrinsic bloat low, one can determine the energy-optimal iteration time and precisely slow down computations in the non-straggler keeping that iteration time (Figure 2(b)).
Figure 2. Two data parallel pipelines run, and the first one became a straggler (e.g., due to the second GPU temporarily overheating). The non-straggler pipeline wastes energy by running as fast as possible. Such extrinsic energy bloat can be removed by precisely slowing down non-straggler pipelines.
Stragglers arise from numerous sources. Thermal or power throttling in a datacenter can result in 10–50% slowdown (A100powerthermal-energies21, ; gpupower-cal23, ; thunderbolt-osdi20, ; mvpp-asplos20, ; polca-arxiv23, ), and I/O bottlenecks in the storage or network can be longer than GPU computation by up to 4×4$\times$4 × (datastalls-vldb21, ; recsys-dsi-isca22, ), acting like a persistent straggler pipeline. Recent failure-resilient training frameworks (bamboo-nsdi23, ; oobleck-sosp23, ) deploy heterogeneous pipelines, introducing non-uniform iteration times. With increasing job and infrastructure scale, the probability of encountering stragglers increases (largescalefailures-sc17, ; philly-atc19, ; mlaas-nsdi22, ; megascale-nsdi24, ; tpuv4-nsdi24, ).
Stragglers are often times known to and anticipated by the training infrastructure, because they were created by the infrastructure itself (e.g., throttling, non-compute bottlenecks, fault-tolerant planning). Therefore, in this work, we focus on how to plan time and energy across time and allow quick adaptation assuming that such information is available.
@@ 2.4. Potential Benefits of Reducing Energy Bloat
To gauge potential energy savings, we measure the energy savings of slowing down every computation in the pipeline to their minimum-energy frequencies. This will slow down the pipeline’s iteration time, but can act as an upper bound for energy savings. For our experimental workloads in Section 6.2, this gives on average 16% and 27% energy reduction on A100 and A40 GPUs, respectively. We show in Section 6.2 that Perseus can realize most of the potential savings with negligible slowdown, even through our problem is NP-hard.
Figure 3. Three cases of where the straggler pipeline’s iteration time $T^{\prime}$ can be. $T_{\min}$ and $T^{*}$ are the shortest and longest iteration time on the “iteration time–energy” Pareto frontier. The black dot is when all computations run at the maximum speed, which wastes energy. The green dot is the energy-optimal iteration time of the non-straggler pipeline. Solid orange dots make up the “iteration time–energy” Pareto frontier, and the orange dotted line shows that energy increases beyond $T^{*}$.
@@ 3. Perseus Overview
Perseus is an energy optimization system for large model training. In this section, we present our unified optimization framework that aims to remove both types of energy bloat (§3.1) and walk through the workflow of Perseus (§3.2).
@@ 3.1. Unified Optimization Framework
Our goal is to minimize a pipeline’s energy consumption by controlling the execution speed of each computation in the pipeline. In doing so, we can slow down a pipeline’s iteration time at most until the straggler’s iteration time $T^{\prime}$:
@@
where $F$ is the set of GPU frequencies [^2] The SM frequency of NVIDIA GPUs can be set via NVML (nvml, ) in around 10 ms (AMD GPUs via AMDSMI (amdsmi, )), which is much shorter than typical large model computation latencies. Locking the GPU’s frequency provides deterministic computation latency (nvml, ; clockwork-osdi20, ), making it suitable for tightly planning and packing execution over time. to run each forward and backward computation in the pipeline, and $\textrm{Time}(F)$ and $\textrm{Energy}(F)$ are the iteration time and energy consumption of the pipeline when executed with $F$, respectively. Changing $F$ will lead to different values of $\textrm{Time}(F)$ and $\textrm{Energy}(F)$ on the “iteration time–energy” 2D plane. However, we are only interested in ($\textrm{Time}(F)$, $\textrm{Energy}(F)$) points that are on the “iteration time–energy” Pareto frontier.
Now, let us assume we have a fully characterized “iteration time–energy” Pareto frontier with $T_{\min}$ and $T^{*}$ bookending the frontier (described in Section 4). $T_{\min}$ is the shortest iteration time on the frontier, which is the same as the iteration time of running every computation at the maximum speed, and $T^{*}$ is the iteration time with minimum energy consumption. Figure 3 depicts the three cases where the straggler’s iteration time $T^{\prime}$ can possibly be w.r.t. the Pareto frontier:
@@
    When there are no stragglers (Figure 3(a)), we simply operate on the point on the Pareto frontier with iteration time $T_{\min}$, which reduces intrinsic energy bloat.
@@
    When a moderately slow straggler is detected (Figure 3(b)), we additionally reduce extrinsic energy bloat while keeping intrinsic bloat low by slowing down all non-straggler pipelines until $T^{\prime}$, using up all the slack time.
@@
    Finally, the straggler’s iteration time may go beyond the minimum-energy point $T^{*}$ on the frontier (Figure 3(c)). In this case, we only slow down non-stragglers until $T^{*}$, because going past $T^{*}$ will instead increase energy.
@@
Therefore, if we pre-characterize the entire “iteration time–energy” Pareto frontier by finding all Pareto-optimal $F$s, Perseus will be able to react instantly to emerging stragglers by looking up $F_{\textrm{opt}}$ that leads to iteration time $T_{\textrm{opt}}$.
Finally, we note that unlike other problem settings that do not consider energy consumption, fully utilizing all the slack time created by the straggler is not always energy-optimal. Being too fast or too slow can both waste energy.

### v3

@@ Abstract.
Training large AI models on numerous GPUs consumes a massive amount of energy, making power delivery one of the largest limiting factors in building and operating datacenters for AI workloads. However, we observe that not all energy consumed during training directly contributes to end-to-end throughput; a significant portion can be removed without slowing down training. We call this portion energy bloat.
In this work, we identify two independent sources of energy bloat in large model training and propose Perseus, a training system that mitigates both. To do this, Perseus obtains the time–energy tradeoff frontier of a large model training job using an efficient graph cut-based algorithm, and schedules computation energy consumption across time to reduce both types of energy bloat. Evaluation on large models, including GPT-3 and Bloom, shows that Perseus reduces the energy consumption of large model training by up to 30% without any throughput loss or hardware modification. [^1] Perseus is open-source as part of Zeus (zeus-nsdi23, ) at https://ml.energy/zeus.
Energy-efficiency, datacenter power management, straggler, distributed training, large model training
journalyear: 2024copyright: acmlicensedconference: ACM SIGOPS 30th Symposium on Operating Systems Principles; November 4–6, 2024; Austin, TX, USAbooktitle: ACM SIGOPS 30th Symposium on Operating Systems Principles (SOSP ’24), November 4–6, 2024, Austin, TX, USAdoi: 10.1145/3694715.3695970isbn: 979-8-4007-1251-7/24/11ccs: Computer systems organizationccs: Computing methodologies Machine learningccs: Software and its engineering Power management
@@ 1. Introduction
As deep neural networks (DNNs) continue to grow in model and dataset size (scalinglaws-arxiv20, ; chinchilla-neurips22, ), the energy consumption of large model training is increasing as well. For instance, training GPT-3 (gpt3, ) reportedly consumed 1.3 GWh (patterson2021carbon, ). Then, this was dwarfed by Amazon’s training of a 200B model, which consumed about 11.9 GWh (hamilton2024constraint, )—enough to power more than 1,000 average US households for a year (us-household, ). Such energy-intensive large model training not only inflates datacenter operational expenses, but also made power delivery a primary challenge in building datacenters today (cbre2023, ; mckinsey2023, ; cbre2024, ; aienergy-joule24, ; openai-keynote-hotchips24, ).
@@
Despite recent works on accelerating large model training (megatronlm-sc21, ; alpa-osdi22, ; merak-tpds23, ), energy optimization remains an open challenge (greenai-cacm20, ; patterson2021carbon, ). While energy optimization is well-studied in the hardware community (eyeriss-jssc16, ; vivienne17efficient, ; puma-asplos19, ; kws-jssc24, ), the power bottleneck of recent datacenters (cbre2023, ; mckinsey2023, ; cbre2024, ; dvfs-boosting-asplos24, ; openai-keynote-hotchips24, ) shows that efficiency gains from hardware advancement alone are not sufficient to sustain the growing demand for AI compute. In light of this, recent works show that software can play a significant role in energy optimization by capturing application characteristics that general-purpose hardware cannot (e.g., no need to finish computation before the deadline), bringing hardware-agnostic energy-efficiency gains (gpoeo-tpds21, ; zeus-nsdi23, ; chase-ccai23, ; envpipe-atc23, ; crosslayer-energy-eecs24, ).
@@
In this paper, we seek a software method that reduces the energy consumption of large model training without slowdown, thereby also reducing average power draw. To that end, we identify energy bloat, the portion of energy consumption that can be removed without slowdown in software systems for large model training. We find two independent sources of energy bloat—intrinsic and extrinsic—and propose a single optimization framework that minimizes both.
@@
Intrinsic energy bloat comes from computation imbalance when a large model is distributed across multiple GPUs with pipeline parallelism (§2.2). Balancing the amount of computation in each pipeline stage is an important problem for distributed execution planning (pipedream-sosp19, ; dapple-ppopp21, ; autopipe-icml21, ; alpa-osdi22, ), but perfectly balancing every stage is not always possible because layers in a DNN are coarse-grained tensor operations with varying amounts of computation. When stages have unequal computation times, those not on the critical path of computation run needlessly fast—that is, they consume energy that does not contribute to the overall training throughput. Such intrinsic energy bloat opens up the opportunity to precisely slow down each non-critical computation in the pipeline such that the length of the critical path does not change.
@@
Extrinsic energy bloat, in contrast, arises when multiple pipelines run in parallel in a synchronous fashion, and one or more pipelines run slower than the rest (§2.3). Root causes behind such slowdowns are varied, including power/thermal throttling (A100powerthermal-energies21, ; gpupower-cal23, ; thunderbolt-osdi20, ; mvpp-asplos20, ; polca-asplos24, ), I/O bottlenecks in the storage/network (datastalls-vldb21, ; recsys-dsi-isca22, ; antdt-icde24, ), and hardware/software failures (bamboo-nsdi23, ; oobleck-sosp23, ; recycle-sosp24, ), and the likelihood of their presence increases with the scale and duration of training (largescalefailures-sc17, ; philly-atc19, ; mlaas-nsdi22, ). All pipelines running faster than the straggler pipeline are needlessly fast, wasting energy that does not affect the overall training throughput. Thus, we can slow down entire pipelines without delaying gradient synchronization.
@@
In this work, we propose Perseus, which formulates a unified optimization framework to remove both intrinsic and extrinsic energy bloat from large model training (§3). At its core, Perseus efficiently pre-characterizes the entire time–energy tradeoff frontier of a training iteration, allowing it to minimize intrinsic bloat under normal operation and to mitigate extrinsic bloat arising from stragglers. Existing works fall short on both fronts. EnvPipe (envpipe-atc23, ) is limited to intrinsic bloat reduction with a point solution that leads to suboptimal energy reduction. Zeus (zeus-nsdi23, ), in contrast, ignores intrinsic bloat as it only considers single-GPU training, which also renders its time–energy frontier suboptimal for large models.
@@
We show that characterizing the optimal time–energy Pareto frontier is NP-hard not only to solve, but also to approximate within any constant factor. Given this impasse, we propose an efficient algorithm that optimally solves a relaxed problem instead (§4). To do so, Perseus represents one training iteration as a directed acyclic graph (DAG) of forward and backward computations in each pipeline stage. Then, Perseus efficiently generates all energy schedules, defined as the planned time and energy consumption of each computation, that are on the time–energy frontier using a graph cut-based algorithm that iteratively crawls up the frontier from the bottom. Minimizing intrinsic and/or extrinsic energy bloat is then as simple as choosing the appropriate energy schedule from the pre-characterized time–energy frontier.
@@
Perseus consists of a client library and a server (§5). The client library integrates with a large model training framework and accelerator to measure computation energy consumption and control accelerator speed. The server produces optimized energy schedules, using the abstract computation DAG and time/energy measurements provided by the client.
@@
Evaluation on large models (GPT-3 (gpt3, ), BERT (bert, ), T5 (t5-jmlr20, ), Bloom (bloom-arxiv22, ), and a scaled-up version of Wide-ResNet (wideresnet-bmvc16, )), shows that Perseus is able to reduce per-iteration energy consumption by up to 30% with negligible or no slowdown, reducing energy consumption and average power draw (§6).
@@
  • We propose Perseus, a software-only energy optimization system that reduces energy bloat through a unified optimization framework and a graph cut-based algorithm.
@@ 2. Motivation
Figure 1. One training iteration of GPT-3 1.3B with 4 pipeline stages and 6 microbatches on NVIDIA A100 GPUs, drawn to scale. For example, F5 and B5 in the S2 row denote forward and backward for the fifth microbatch on Stage 2. The critical path is traced with a blue line. Colors show power consumption. Other models are visualized in Appendix A.
@@ 2.1. Large Model Training
Large model training is mostly dominated by 3D (data, tensor, and pipeline) parallelism (megatronlm-sc21, ; megatron-turing-arxiv22, ; bloom-arxiv22, ; merak-tpds23, ; falcon-arxiv23, ). Especially, pipeline parallelism partitions a large model into multiple stages and its training batch into microbatches, and pipelines forward and backward computations through the stages. Then, such pipelines are replicated to perform data parallel training. Pipelines can only move on to the next iteration after every pipeline has finished and synchronized gradients.
@@ 2.2. Intrinsic Energy Bloat
We profile GPT-3 1.3B on NVIDIA A100 GPUs and visualize the timeline of one training iteration in Figure 1(a). In addition to the familiar bubbles in the 1F1B schedule (megatronlm-sc21, ), we observe gaps between forward and backward computations, where the GPU is simply blocking on communication with an adjacent stage. Such gaps exist because the computation time of each pipeline stage is not perfectly balanced. Partitioning stages in a balanced manner is an important problem in distributed execution planning (gpipe-neurips19, ; pipedream-sosp19, ; dapple-ppopp21, ; alpa-osdi22, ), but perfect balancing is difficult because DNNs are essentially a sequence of coarse-grained tensor operations with varying sizes.
@@
Table 1. Forward latency ratio of the longest to the shortest stage on A100 GPUs. 1.00 would mean perfect balance.
@@
Given stage imbalance, not all forward and backward computations are on the critical path of computation (Figure 1(a)). This means that non-critical computations running at their maximum speed are not contributing to faster iteration time, and thus simply wasting energy. We call this intrinsic energy bloat, which can be reduced by precisely slowing down each non-critical computation without lengthening the critical path (Figure 1(b)). Although seemingly simple, this problem is not only NP-hard to solve, but also NP-hard to even approximate to any constant factor (svensson2012hardness, ).
@@ 2.3. Extrinsic Energy Bloat
Numerous replicas of the same pipeline run in a data parallel fashion in large model training. Because every pipeline must synchronize gradients at the end, if one pipeline runs slower, all other pipelines must wait until the straggler pipeline finishes (Figure 2(a)). Since the straggler pipeline determines end-to-end iteration time, all other pipelines running at their fastest possible iteration time are wasteful. We call this extrinsic energy bloat, because unlike intrinsic energy bloat, its cause is extrinsic to the training pipeline. To reduce extrinsic bloat while keeping intrinsic bloat low, one can determine the energy-optimal iteration time for non-straggler pipelines and precisely slow down computations so that non-straggler pipelines attain that iteration time (Figure 2(b)).
Figure 2. Among two data parallel pipelines, the first one becomes a straggler. The non-straggler pipeline causes extrinsic energy bloat by running as fast as possible (a), which can be reduced by precisely slowing it down (b).
Stragglers arise from numerous sources. Thermal or power throttling in a datacenter can result in 10–50% slowdown (A100powerthermal-energies21, ; gpupower-cal23, ; thunderbolt-osdi20, ; mvpp-asplos20, ; polca-asplos24, ), and I/O bottlenecks in the storage or network can be longer than GPU computation by up to 4×4$\times$4 × (datastalls-vldb21, ; recsys-dsi-isca22, ; antdt-icde24, ), acting like a persistent straggler pipeline. Recent failure-resilient training frameworks (bamboo-nsdi23, ; oobleck-sosp23, ; recycle-sosp24, ) deploy heterogeneous pipelines, introducing non-uniform iteration times. With increasing job and infrastructure scale, the probability of encountering stragglers increases (largescalefailures-sc17, ; philly-atc19, ; mlaas-nsdi22, ; megascale-nsdi24, ; tpuv4-nsdi24, ).
In this work, we focus on stragglers that are known to and anticipated by the training infrastructure, generally because they were created by the infrastructure itself (e.g., power and thermal throttling, non-compute bottlenecks, fault-tolerant planning). Such stragglers also tend to persist beyond typical training iteration times. Therefore, Perseus focuses on planning time and energy consumption across time and allowing quick adaptation, assuming that information about stragglers is available.
@@ 2.4. Potential Benefits of Reducing Energy Bloat
To gauge potential energy savings, we measure the energy savings achieved by slowing down every computation in the pipeline to their minimum-energy frequencies. This will slow down iteration time, but can act as an upper bound for energy savings. For our workloads in Section 6.2, this gives on average 16% and 27% energy reduction on A100 and A40 GPUs, respectively. Section 6.2 shows that Perseus can realize most of the potential savings with negligible slowdown.
Figure 3. Three cases showing where the straggler pipeline’s iteration time $T^{\prime}$ can be. $T_{\min}$ and $T^{*}$ are the shortest and longest iteration times on the time–energy frontier. The black dot is when all computations run at the maximum speed, which wastes energy. The green dot is the energy-optimal iteration time of the non-straggler pipeline. Solid orange dots make up the frontier, and the orange dotted line shows that iteration energy increases beyond $T^{*}$.
@@ 3. Perseus Overview
We first present Perseus’s unified optimization framework that aims to remove both types of energy bloat (§3.1), and then walk through the workflow of Perseus (§3.2).
@@ 3.1. Unified Optimization Framework
Our goal is to minimize a pipeline’s energy consumption by controlling the execution speed of each computation in the pipeline. In doing so, we can slow down a pipeline’s iteration time up to the straggler’s iteration time $T^{\prime}$:
@@
where $F$ is an assignment of GPU frequencies [^3] The SM frequency of NVIDIA GPUs can be set via NVML (nvml, ) in around 10 ms, which is much shorter than typical large model computation latencies. Locking the GPU’s frequency provides deterministic computation latency (nvml, ; clockwork-osdi20, ), making it suitable for tightly planning and packing execution over time. to each forward and backward computation in the pipeline, and $\textrm{Time}(F)$ and $\textrm{Energy}(F)$ are the iteration time and energy consumption of the pipeline when executed with $F$, respectively. Changing $F$ will lead to different values of $\textrm{Time}(F)$ and $\textrm{Energy}(F)$, but we are only interested in ($\textrm{Time}(F)$, $\textrm{Energy}(F)$) points that are on the time–energy tradeoff frontier.
Now, let us assume we have a fully characterized time–energy frontier, bookended by $T_{\min}$ and $T^{*}$ (Section 4 is dedicated to describing how). $T_{\min}$ is the shortest iteration time on the frontier, which is the same as the iteration time of running every computation at the maximum speed, and $T^{*}$ is the iteration time with minimum energy consumption, which is when each computation runs at the frequency that consumes the least amount of energy for that computation. [^4] This is typically not the lowest frequency, because computations running with very low frequencies incur more latency increase than power reduction, resulting in higher energy consumption. Figure 3 shows the three possible cases regarding where the straggler’s iteration time $T^{\prime}$ can be:
@@
    Figure 3(a): When there are no stragglers, we simply select the point on the frontier with iteration time $T_{\min}$, which reduces only intrinsic energy bloat.
@@
    Figure 3(b): When a moderately slow straggler is detected, we additionally reduce extrinsic energy bloat while keeping intrinsic bloat low by slowing down all non-straggler pipelines until $T^{\prime}$, using up all the slack time.
@@
    Figure 3(c): Finally, the straggler’s iteration time may go beyond the minimum-energy point $T^{*}$ on the frontier. In this case, we only slow down non-stragglers until $T^{*}$, because going past $T^{*}$ will instead increase energy.
@@
Therefore, when a straggler emerges (i.e., $T_{\min}<T^{\prime}$), Perseus can compute $T_{\textrm{opt}}$ using Equation 2 and quickly look up the frequency plan $F_{\textrm{opt}}$ that leads to iteration time $T_{\textrm{opt}}$ using the pre-characterized time–energy frontier.
Finally, we note that, unlike other problem settings that do not consider energy consumption, fully utilizing all the slack time created by the straggler is not always energy-optimal; being too fast or too slow can both waste energy.